OpenStudy (anonymous):

Two quantities vary inversely. If the value of the first is 15 when the value of the second is 18, find the value of the second quantity when the first is 10.

OpenStudy (wasiqss):

24

OpenStudy (anonymous):

p varies directly as q. When q = 31.2, p = 20.8. Find p when q = 15.3. and that one :)

OpenStudy (lgbasallote):

y = k/x will be your equation since it's inversely proportional 15 = k/18 270 = k 10 = 270/x 10x = 270 x = 27

OpenStudy (anonymous):

x=k/y 15=k/18 k=270 for x=10, 10=270/y y=27 the 2nd number is 27

OpenStudy (anonymous):

LGBAAAA wins :D

OpenStudy (lgbasallote):

wasiqss please do not post answers only..it is against the revised code of conduct..the mods would be mad..we dont want that to happen...

OpenStudy (anonymous):

haha can you guys help wit hteh second one too

OpenStudy (wasiqss):

lgb duhhh

OpenStudy (lgbasallote):

p = kq 20.8 = k 31.2 k = 20.8/31.2 k = 0.66 p = 0.66 (15.3) p = 10.2

OpenStudy (anonymous):

@Was be nice

OpenStudy (wasiqss):

ok